OpenStudy (anonymous):

Daniel set a goal for the number of servings of fruits and vegetables he wants to eat each day. He has already eaten two and one-fourth servings today, and he still needs to eat three and three-fourths servings to meet his goal. What is his goal for daily servings of fruits and vegetables?

OpenStudy (anonymous):

2 1/4 + 3 3/4 =______ number of servings

OpenStudy (anonymous):

2+3 =5 1/4 + 3/4 = 1 5+1 = 6 answer is 6